# Workload Estimation for Unknown Tasks: A Survey of Machine Learning Under Distribution Shift

### v1

@@ Abstract
Human-robot teams involve humans and robots collaborating to achieve tasks under various environmental conditions. Successful teaming will require robots to adapt autonomously to a human teammate’s internal state. An important element of such adaptation is the ability to estimate the human teammates’ workload in unknown situations. Existing workload models use machine learning to model the relationships between physiological metrics and workload; however, these methods are susceptible to individual differences and are heavily influenced by other factors. These methods cannot generalize to unknown tasks, as they rely on standard machine learning approaches that assume data consists of independent and identically distributed (IID) samples. This assumption does not necessarily hold for estimating workload for new tasks. A survey of non-IID machine learning techniques is presented, where commonly used techniques are evaluated using three criteria: portability, model complexity, and adaptability. These criteria are used to argue which techniques are most applicable for estimating workload for unknown tasks in dynamic, real-time environments.
Index Terms:
@@
human-robot teams, non-iid machine learning, workload estimation for unknown tasks
publicationid: pubid:
I Introduction
@@
Deploying human-robot teams in uncertain environments requires robots to have a dynamic understanding of humans’ internal state. This dynamic understanding must account for real world complexities in order to enable fluid interactions between a robot and a human teammate. Workload represents how hard a person is working, and can be decomposed into workload components (e.g., cognitive, speech, auditory, visual, gross motor, fine motor, and tactile). Incorporating accurate workload models will enhance the robot’s understanding of its human teammate, allowing the robot to manage a human’s workload level, either through intelligent task redistribution or modulation of how the robot interacts with the human. However, real world human-robot teams will inevitably encounter unseen tasks and environments.
@@
An adaptive teaming system requires machine learning models that can estimate workload accurately in both known and unknown situations. Further, workload estimates are susceptible to individual differences across individuals performing the same task, the same individual performing different tasks, and even for the same individual performing the same task on different days. Workload is also influenced by factors, such as experience, stress, and fatigue [1]. Generalizing to novel situations while also accounting for these factors is difficult for traditional machine learning models, as these standard approaches assume data consists of independent and identically distributed (IID) samples [2]. However, this assumption does not hold for estimating workload for new tasks.
@@
Consider first response robotics. Urban firefighters will work alongside robots to perform a wide range of tasks [3, 4]. Firefighters have historically used robots to gather aerial imagery of buildings, but more advanced control algorithms will enable the use of similar robots to extinguish fires in difficult to access locations [4]. This new tasks require different cognitive and visual demands, as well as different degrees of robot control, supervision, and interactions.
Prior work established that standard machine learning methods are capable of generalizing across individuals for similar tasks (e.g., [1, 5, 6]), but there is no widely accepted approach that can generalize to unknown tasks [7]. Generalizing across tasks is difficult primarily due to a shift in the relative balance of workload components; thus, various physiological signals are of particular importance to a specific task and the magnitude of that significance varies between situations [1].
This dynamic relationship between physiological signals and an individual’s underlying workload is exemplar of the variability of human behavior, which makes the development of machine learning models that generalize in the real world difficult. Gathering the necessary data to develop these models exacerbates this problem, as experimental human-subject evaluations rarely capture the fully breadth of human activity within real world domains. Similar complexity is seen across human-robot interaction (HRI) application domains [8, 9, 10]. The real world variability of human behavior and activity violates the IID assumption, requiring non-standard machine learning techniques (e.g., continual learning, meta-learning) to successfully overcome these challenges [11].
The manuscript surveys applicable non-IID machine learning methods. Specifically, the non-IID methods are evaluated to assess their viability for developing models of the human’s internal state (e.g., workload estimation) that can accommodate real world variability. Workload estimation for unknown tasks is an exemplar of the challenge in constructing machine learning models for HRI application domains that are characterized by non-IID data. Section II discusses the fundamental of objective workload estimation and distribution shifts that characterize non-IID machine learning. Section III lays out the notation and evaluation criteria. Section IV evaluates a wide range of non-IID machine learning methods, and Section V discusses the most applicable methods for workload estimation for unknown tasks. Finally, Section VI concludes and presents potential future work.
II Background
@@
Workload is a complex, dynamic, individual-specific, non-linear construct, and developing models that can account for these complexities is a non-trivial task. Fundamentally, estimating workload for unknown tasks is a problem of generalization and recent work has established that existing workload assessment methods encounter difficulties generalizing across tasks, either known or unknown [7] [12]. Modern techniques (e.g., meta-learning) approach machine learning from a different perspective and possess the potential to build more flexible workload estimation models.
@@
Humans have limited resources for processing information, making decisions, and dealing with physical stress. Formally, workload is the degree of activation of a finite pool of resources while performing a task [7] that is influenced by environmental factors, situational factors, and internal characteristics (e.g., fatigue, experience). Workload has been shown to vary across individuals, across tasks, and over time [13] [14]. Workload can be divided into different components based upon the types of resources being utilized: cognitive, speech, auditory, visual, and physical [15], and physical can be further subdivided based upon the nature of physical work: gross motor, fine motor, and tactile [16].
@@
Workload estimation methods either rely on subjective (e.g., NASA Task Load Index [17], In-Situ Surveys [18]) or objective metrics. Subjective metrics fail to provide a continuous measure and are sensitive to the bias inherent to self-reports [19] [20]. Robots operating in dynamic environments need high-frequency workload measures to adequately adjust to their human teammates’ current state. Objective estimation methods use machine learning to learn the relationship between physiological metrics, collected via wearable sensors, and an individual’s underlying workload (e.g., [21, 22, 23, 24, 25]).
Human-robot teams that are deployed in unstructured, dynamic environments cannot rely on environmentally embedded sensors (e.g., built-in cameras), or complex sensor systems (e.g., electroencephalogram (EEG)). Thus, wearable sensors (e.g., a heart-rate monitor) can serve as the primary data source, enabling the human and the robot to act independently. Each workload component corresponds to a different set of physiological metrics, though some metrics may correspond to multiple components. A good example is pupil diameter, which is positively correlated with cognitive and visual workload [16]. Robots need a reliable estimate of the distribution of workload across the components, as tasks will change over time, and differing tasks impact different components.
@@
Prior work established four metrics for evaluating workload estimation algorithms: Sensitivity, Diagnosticity, Suitability, and Generalizability [16]. Sensitivity represents an algorithm’s ability to reliably detect workload levels (i.e., $\geq$ 80% accuracy or $\leq$ 5 root mean squared error (RMSE)). Diagnosticity refers to the capability of detecting different types of workload. Suitability refers to an algorithm’s ability to assess the complete overall workload state. The last metric, generalizability, refers to the algorithm’s ability to generalize across individuals and tasks. This metric defines “generalizing across tasks” as an algorithms capacity to asses the complete overall workload state (i.e., each workload component) and defines “generalizing across individuals” as an algorithms capacity to maintain an accuracy of $\geq$ 80% or an RMSE $\leq$5 for all individuals.
@@
Most existing methods fail to meet these criteria, as they primarily perform binary classification of cognitive workload (e.g., [26, 27, 28, 29]). Many methods are either suitable (e.g., [30, 31, 32, 33]) or sensitive (e.g., [34, 35, 5, 36]), but very few algorithms are diagnostic. Further, little work has been done to evaluate the generalizability of existing workload estimation methods.
@@
Prior work tends to focus on either generalizing across individuals or generalizing across tasks. Many algorithms employed a leave-one-subject-out cross-validation strategy to validate generalization across individuals (e.g., [37, 24, 38, 39]); however, generalizing across tasks, either known or unknown, is rarely evaluated. Algorithms specifically developed to generalize across tasks have received increased attention, and are typically referred to as cross-task workload estimation algorithms [40]. Much of this work evaluated an existing algorithm’s ability to generalize across tasks, but those algorithms did not achieve either a $\geq 80\%$ accuracy, or a $\leq$ 5 RMSE on arbitrary tasks (e.g., [40, 41, 42, 43, 44]). Thus, these algorithms are sensitive, but not generalizable. Other algorithms achieved $\geq 80\%$ accuracy across tasks, but only perform binary classification (e.g., [45, 46, 47]); thus, failing to meet the generalizability criterion.
@@
A support vector machine-based workload estimation algorithm that uses Recursive Feature Elimination to optimize the input features was shown to estimate cognitive workload on arbitrary tasks [48]. This method achieved $<$ 1.0 mean squared error on cross-task workload estimation. Similar methods were developed by Guan et al. [23] and Zhou et al. [49]. Additionally, deep learning techniques (e.g., attention mechanisms) that estimate workload have shown promise [50]. However, the majority of existing methods focus solely on cognitive workload and rely on high data volumes from a static sensor (i.e., EEG [51, 52, 53]). Thus, none of the methods are viable for use in dynamic, uncertain environments.
@@
Real world human-robot teams will inevitably encounter unique scenarios when deployed. An adaptive teaming system requires models that can accurately estimate workload in both known and unknown scenarios, as well as accommodate individual differences [14]. Generalizing to unknown tasks while also accounting for individual differences is difficult for traditional machine learning, as the standard approaches assume data consists of independent and identically distributed (IID) samples [2]. Generalizing across tasks is difficult primarily due to a shift in the relative balance of the workload components associated with a task; thus, various physiological signals are of particular importance to a specific task and the magnitude of that significance varies between situations [1]. This shift in the balance of workload components and physiological signals violates the IID assumption, which requires methods that use non-standard machine learning techniques [2].
III Notation and Evaluation Criteria
Dynamic human-robot teams will rely on machine learning methods to enable a robot to quantify different aspects of their human teammate (e.g., workload). A major challenge is that these teams will continuously encounter new scenarios that the robot’s underlying machine learning methods cannot account for. These new scenarios constitute a meaningful distribution shift from conventional machine learning method’s training dataset (i.e., violate the IID assumption). Specifically, a distribution shift indicates scenarios where training and testing data come from different distributions. Understanding the difference between non-IID machine learning methods is key for building human-robot teams that can handle the variability of real world problem domains.
@@ III Notation and Evaluation Criteria
The workload estimation and the machine learning literature both cover a large spectrum of work. Naturally, there are terms that frequently occur in both fields which require disambiguation. The term task is used in the workload literature to refer to a job an individual may perform (e.g., lifting boxes, mental calculations), but the machine learning literature uses the term to refer to a learning task for a machine learning model. The term domain is often used to characterize the input space for a machine learning problem. For example, performing object detection under various weather conditions (e.g., sunny, cloudy) can be viewed as two domains characterized by the weather [54]. Domain may also refer to a set of activities a human performs for a particular job (e.g., wildland firefighters).
@@
The term learning task will refer to instances of a machine learning problem. The term task, without any adjectives, will refer to a task conducted by a human. The term domain will refer to different input spaces that a machine learning algorithm may operate on. The term application domain will refer to applications where human-robot teams may be deployed.
@@
Definitions of domain and learning task, in the context of distribution shift, have long been established [55]:
Definition III.1
@@
(Domain) A domain $D$ is a two-tuple $(\mathcal{X},P(X))$. $\mathcal{X}$ is the feature space of $D$ and $P(X)$ is the marginal distribution, where $X=\{x_{1},…,x_{n}\}\in\mathcal{X}$.
Definition III.2
@@
Grounding this notation in workload estimation helps to provide an intuitive explanation of the different variables. A particular workload estimation instance represents the domain, $D$, which is characterized by the set of physiological metrics, $\mathcal{X}$. The marginal distribution, $P(X)$, represents the probability distribution over the physiological metrics’ possible values.
@@
A typical task recognition problem is defined as training a predictive function to learn the relationship between a set of physiological metrics and a label space (e.g., (1, 2, and 3)), where each integer represents a unique task. A similar task recognition problem can be defined using the same metrics and the same label space, but where each integer represents the task difficulty (i.e., easy, medium, or hard). These two task recognition problems are semantically different, but the underlying machine learning method does not inherently recognizes these differences; therefore, uniqueness of these two learning tasks cannot be fully specified by the domain and label space. The changes in how the resulting predictive functions, $f(x)$, map the physiological metrics to the output values is a key characteristic in differentiating these two learning tasks.
Numerous methods have been applied to the non-IID machine learning problems, each characterized by the number of distributions and distribution shift type (see Fig. 1). Three evaluation criteria are applied to the presented methods: (1) Portability (2) Model Complexity (3) Adaptability.

### v2

# A Survey of Machine Learning for Estimating Workload: Considering Unknown Tasks

@@ Abstract
Successful human-robot teaming will require robots to adapt autonomously to a human teammate’s internal state, where a critical element of such adaptation is the ability to estimate the human’s workload in unknown situations. Existing workload models use machine learning to model the relationship between physiological signals and workload. These methods often struggle to generalize to unknown tasks, as the relative importance of various physiological signals change significantly between tasks. Many of these changes constitute a meaningful shift in the data’s distribution, which violates a core assumption made by the underlying machine learning approach. A survey of machine learning techniques designed to overcome these challenges is presented, where common techniques are evaluated using three criteria: portability, model complexity, and adaptability. These criteria are used to analyze each technique’s applicability to estimating workload during unknown tasks in dynamic environments and guide future empirical experimentation.
keywords:
@@
human-robot teams, workload estimation for unknown tasks, machine learning under distribution shift
1 Introduction
@@
Deploying human-robot teams in uncertain environments requires robots to have a dynamic understanding of humans’ internal state. This dynamic understanding must account for real-world complexities in order to enable fluid interactions between a robot and a human teammate. Workload represents how hard a person is working, and can be decomposed into workload components (i.e., cognitive, speech, auditory, visual, gross motor, fine motor, and tactile).
Incorporating machine learning-based workload models that accurately capture the contribution of each component will enhance the robot’s understanding of its human teammate, allowing it to respond appropriately to undesirable workload levels (i.e., overload (OL) or underload (UL)) through intelligent modulation of how the robot interacts with the human. These adaptive interactions enable more collaborative human-robot teaming dynamics and facilitate long-term collaboration. However, real-world human-robot teams will inevitably encounter novel (i.e., unknown) tasks. These tasks present unique challenges for the machine learning models that underpin modern workload estimation algorithms that rely on physiological signals, as unknown tasks frequently exhibit unique data characteristics that the model was not trained to account for.
@@
An adaptive teaming system requires machine learning models that can estimate workload accurately in both known and unknown situations. Further, workload estimates are susceptible to individual differences across individuals performing the same task, the same individual performing different tasks, and even for the same individual performing the same task on different days (Longo et al., 2022; Wickens et al., 2004). Workload is also influenced by factors, such as experience, stress, and fatigue (Heard et al., 2019). Developing machine learning-based workload estimation models that generalized to novel situations while also accounting for these factors is difficult, as these traditional approaches assume data consists of independent and identically distributed (IID) samples (Murphy, 2023, chap. 19). The IID assumption states that a machine learning model’s training and testing data (e.g., real-world data) are drawn from the same probability distribution. This assumption implies that the training data is sufficiently representative of real-world data, which does not necessarily hold when estimating workload for novel, unknown tasks.
@@
Prior work established that workload models built using standard machine learning methods are capable of generalizing across individuals for similar tasks (e.g., Heard et al., 2019; Kaczorowska et al., 2021; Manawadu et al., 2018), but there is no widely accepted approach that can generalize to novel, unknown tasks (Longo et al., 2022). Generalizing across tasks is difficult due to a shift in the relative balance of workload components, where certain physiological signals and workload components are of particular importance to a given task and the magnitude of that significance varies between tasks (Heard et al., 2019). Consider first response robotics. Urban firefighters will work alongside robots to perform many tasks (Delmerico et al., 2019). Firefighters have historically used robots to gather aerial imagery, but advanced control algorithms will enable robots to extinguish fires in otherwise inaccessible locations (Perez-Saura et al., 2023). This new task will require different cognitive and visual demands, as well as different degrees of robot control, supervision, and interactions.
The dynamic relationship between physiological signals and an individual’s underlying workload makes the development of machine learning models that generalize in the real world difficult (Albuquerque et al., 2022; Popivanov and Mineva, 1999). Prior work demonstrated that this dynamism is considerably influenced by contextual factors (e.g., tasks demands, teaming dynamics), as contextual factors heavily influence an individual’s workload (Heard et al., 2018; Longo et al., 2022) and their physiological response to external stimuli (Larradet et al., 2020; Saganowski et al., 2022). Considering how contextual factors influence each workload components adds additional complexity. Each component relies on a unique subset of physiological signals (Bhagat Smith et al., 2022, 2023; Fortune et al., 2020; Heard et al., 2019), but it is impossible to predict the specific physiological response novel tasks will incur. Gathering the necessary data to develop workload estimation models exacerbates this problem, as experimental human-subject evaluations rarely capture the fully breadth of real-world human activity (Bhagat Smith et al., 2024).
The variability between contexts (e.g., tasks) violates the IID assumption (Albuquerque et al., 2022; Cao, 2014; Popivanov and Mineva, 1999), and must be analyzed through the lens of distribution shifts (i.e., the mathematical differences between datasets). Prior work developed a broad range of non-IID machine learning techniques (e.g., continual learning, meta-learning) that incorporate contextual examples to ameliorate a distribution shift’s negative impact (Quiñonero-Candela et al., 2022). These techniques have demonstrated success in many application domains (e.g., computer vision (Tian et al., 2020), natural language (Latif et al., 2022), task recognition (Leite and Xiao, 2022)), but their success varies based upon several dataset characteristics (e.g., labeling, volume). There exists too many techniques for a comprehensive evaluation to be tractable; thus, a more careful analysis evaluating how relevant human factors interact with a technique’s machine learning characteristics is required.
The manuscript surveys applicable non-IID machine learning methods. These methods are evaluated to assess their viability for developing workload estimation models that can accommodate real-world variability. First, the fundamentals of objective workload estimation and distribution shifts are presented, followed by requisite notation and criteria. These criteria are used to analyze a wide range of non-IID machine learning techniques, where arguments for which techniques merit empirical investigation are presented. The most applicable techniques are further discussed, which highlights the importance of jointly evaluating a problem’s machine learning and human factors considerations in order to determine a techniques real-world applicability. Finally, conclusions and future work are presented.
2 Background
@@
Workload is a complex, dynamic, individual-specific, non-linear construct, and developing models that can account for these complexities is a non-trivial task. Fundamentally, estimating workload for unknown tasks is a problem of generalization and recent work has established that existing workload assessment methods encounter difficulties generalizing across tasks, either known or unknown (Longo et al., 2022; Zhao et al., 2018).
@@
Humans have limited resources for processing information, making decisions, and dealing with physical stress. Formally, workload is the degree of activation of a finite pool of resources while performing a task (Longo et al., 2022) that is influenced by environmental factors, situational factors, and internal characteristics (e.g., fatigue, experience). Workload has been shown to vary across individuals, across tasks, and over time (Christensen et al., 2012; Wickens et al., 2004). Workload can be divided into different components based upon the types of resources being utilized: cognitive, speech, auditory, visual, and physical (Mitchell, 2000), and physical can be further subdivided based upon the nature of physical work: gross motor, fine motor, and tactile (Heard et al., 2018).
@@
Robots operating in dynamic environments need high-frequency workload measures to adequately adjust to their human teammates’ current state. Workload estimation methods either rely on subjective or objective metrics. Surveys and questionnaires (e.g., NASA Task Load Index (Hart and Staveland, 1988), In-Situ Surveys (Wilson and Russell, 2003)) are common subjective metrics that rely on an individual to self-assess their perceived workload levels (Heard et al., 2018). Subjective metrics fail to provide a continuous measure and are sensitive to the bias inherent to self-reports (Kosch et al., 2023; Matthews et al., 2020); thus, they are not viable for real-world environments. Physiological signals provide a quantitative and consistent method for measuring a human’s workload and are the most widely used objective metric (Debie et al., 2019; Guan et al., 2022). Objective workload estimation methods use machine learning to learn the relationship between physiological metrics, collected via wearable sensors, and an individual’s underlying workload (e.g., Cao et al., 2021; Guo et al., 2021; Quan et al., 2023). These methods provide workload estimates at the necessary frequency and accuracy for real-world environments (Bagheri and Power, 2022; Fortune et al., 2020; Heard et al., 2019).
Many objective workload estimation methods are ill suited for uncertain, dynamic environments as the sensors required to collect the necessary physiological signals are either complex systems (e.g., electroencephalogram (EEG) (Ved and Yildirim, 2021)) or environmentally embedded (e.g., built-in cameras (Kosmopoulos et al., 2012)). The human must be able to move through the environment freely. Environmentally embedded sensors restrict the human’s movement and complex sensor systems are extremely sensitive to the noise introduced by human movement, making them ill suited for uncertain, dynamic environments. Wearable sensors (e.g., a heart-rate monitor) can serve as the primary data source, enabling the human and the robot to act independently. Wearable sensors are susceptible to high sensor noise and latency issues, but prior work has demonstrated their ability in accurately estimating workload in real-time environments (Dayal et al., 2024; Heard et al., 2020; Liu et al., 2024; Park et al., 2024).
Robots need a reliable estimate of the distribution of workload across the workload components, as tasks will change over time, and differing tasks impact different components. Each workload component corresponds to a different set of physiological metrics, though some metrics may correspond to multiple components. Incorporating multiple wearable sensors and physiological signals allows a workload estimation method to differentiate between the workload components and accurately measure their contribution to overall workload (Heard et al., 2018; Bhagat Smith et al., 2024).
@@
Prior work established four metrics for evaluating workload estimation algorithms: Sensitivity, Diagnosticity, Suitability, and Generalizability (Heard et al., 2018). Sensitivity represents an algorithm’s ability to reliably detect workload levels (i.e., $\geq$ 80% accuracy or $\leq$ 5 root mean squared error (RMSE)). Diagnosticity refers to the capability of detecting different types of workload. Suitability refers to an algorithm’s ability to assess the complete overall workload state. The last metric, generalizability, refers to the algorithm’s ability to generalize across individuals and tasks. This metric defines “generalizing across tasks” as an algorithm’s capacity to asses all seven workload components and defines “generalizing across individuals” as an algorithm’s capacity to maintain an accuracy of $\geq$ 80% or an RMSE $\leq$5 for all individuals.
@@
Most existing methods fail to meet these criteria, as they primarily perform binary classification of cognitive workload (e.g., Dell’Agnola et al., 2021; Hogervorst et al., 2014; Xie et al., 2019; Yin et al., 2019). Many methods are either exhibit suitability (e.g., Islam et al., 2020; Momeni et al., 2019; Moustafa et al., 2017; Zhang et al., 2020) or sensitivity (e.g., Caywood et al., 2017; Ding et al., 2020; Kaczorowska et al., 2021; Ved and Yildirim, 2021), but very few exhibit diagnosticity. Further, little work evaluated the generalizability of existing workload estimation methods.
@@
Prior work tends to focus on generalizing across individuals. Many algorithms employed a leave-one-subject-out cross-validation strategy to validate generalization across individuals (e.g., Albuquerque et al., 2019; Guo et al., 2021; Heard et al., 2019; Hefron et al., 2017; Novak et al., 2015); however, generalizing across tasks, either known or unknown, is rarely evaluated. Algorithms specifically developed to generalize across tasks have received increased attention, and are typically referred to as cross-task workload estimation algorithms (Appel et al., 2023). Many of these efforts evaluated an existing algorithm’s ability to generalize across tasks, but those algorithms did not achieve either a $\geq 80\%$ accuracy, or a $\leq$ 5 RMSE on arbitrary tasks (e.g., Appel et al., 2023; Baldwin and Penaranda, 2012; Besson et al., 2012; Boring et al., 2020; Walter et al., 2013). Thus, these algorithms exhibit sensitivity, but not generalizability. Other algorithms achieved $\geq 80\%$ accuracy across tasks, but only perform binary classification (e.g., Dimitrakopoulos et al., 2017; Kakkos et al., 2021; Zhang et al., 2018); thus, failing to meet the generalizability criterion.
@@
Some support vector machine-based workload estimation algorithms have achieved achieved $<$ 1.0 mean squared error on cross-task workload estimation (Guan et al., 2022; Ke et al., 2015; Zhou et al., 2022). Deep learning techniques (e.g., attention mechanisms) that estimate workload have also shown promise (Taori et al., 2022). However, the majority of existing methods focus solely on cognitive workload and rely on high data volumes from complex or environmentally embedded sensors (e.g., Ji et al., 2023; Guan et al., 2023; Wen et al., 2023). Thus, none of the methods are viable for use in dynamic, uncertain environments.
@@
Real-world human-robot teams will inevitably encounter unique tasks when deployed. An adaptive teaming system requires models that can accurately estimate workload in both known and unknown scenarios, as well as accommodate individual differences (Wickens et al., 2004). Generalizing to unknown tasks while also accounting for individual differences is difficult for traditional machine learning, as the standard approaches assume data consists of independent and identically distributed (IID) samples (Murphy, 2023, chap. 19). Generalizing across tasks is difficult primarily due to a shift in the relative balance of the workload components associated with a task; thus, various physiological signals are of particular importance to a specific task and the magnitude of that significance varies between situations (Heard et al., 2019). This shift in the balance of workload components and physiological signals violates the IID assumption, which requires methods that use non-standard machine learning techniques (Murphy, 2023, chap. 19).
3 Notation and Evaluation Criteria
Dynamic human-robot teams will employ robots that estimate their human teammate’s workload using machine learning methods. Extending the robot’s capabilities to accommodate unknown tasks poses unique machine learning challenges, as these tasks constitute a meaningful distribution shift. A distribution shift indicates a scenario where training and testing data come from different probability distributions. Understanding the difference between applicable machine learning methods is key for building human-robot teams that can handle the real-world complexity. The workload estimation and the machine learning literature both cover a large spectrum of work. There are frequently occurring terms in both fields; thus, analyzing applicable methods requires disambiguating key terminology and establishing evaluation criteria to ensure consistent analysis.
3.1 Notation
@@
The term task in the workload literature refers to a job an individual may perform (e.g., lifting boxes), but the machine learning literature uses the term to refer to a learning task for a machine learning model. The term domain is used to characterize the input space for a machine learning problem. For example, performing object detection under various weather conditions (e.g., sunny, cloudy) can be viewed as two domains characterized by the weather (Sun et al., 2022). Domain can also refer to a set of activities a human performs for a particular job (e.g., wildland firefighters).
@@
The term learning task will refer to a machine learning problem. The term task, without adjectives, will refer to an activity conducted by a human. The term domain will refer to different input spaces that a machine learning algorithm may operate on. The term application domain will refer to applications where human-robot teams may be deployed.
@@
The definitions of domain and learning task, in the context of distribution shift, are established (Pan and Yang, 2009):
Definition:.
@@
(Domain) A domain $D$ is a two-tuple $(\mathcal{X},P(X))$. $\mathcal{X}$ is the feature space of $D$ and $P(X)$ is the marginal distribution, where $X=\{x_{1},\dots,x_{n}\}\in\mathcal{X}$.
Definition:.
@@
Grounding this notation in workload estimation provides an explanation of the different variables. A particular workload estimation instance represents the domain, $D$, which is characterized by the set of physiological metrics, $\mathcal{X}$. The marginal distribution, $P(X)$, represents the probability distribution over the physiological metrics’ possible values.
@@
A typical task recognition problem is defined as training a predictive function to learn the relationship between a set of physiological metrics and a label space (e.g., [1, 2, and 3]), where each integer represents a unique task. A similar task recognition problem can be use the same metrics and the same label space, but with each integer representing the task difficulty (i.e., [easy, medium, or hard]). These two problems are semantically different, but the underlying machine learning method does not inherently recognize these differences; therefore, the uniqueness of these two learning tasks cannot be fully specified by the domain and label space. The changes in how the resulting predictive functions, $f(x)$, map the physiological metrics to the output values is a key characteristic in differentiating these two learning tasks.
3.2 Evaluation Criteria
An exhaustive analysis of applicable machine learning methods is impractical due to the volume of existing methods applied to non-IID machine learning problems; however, these methods can be conceptually grouped based upon the number of distributions and the distribution shift type (see Fig. 1) (Murphy, 2023). These categories represent a comprehensive list of non-IID machine learning method types and each category is assessed to understand which method types merit further investigation. Three evaluation criteria are applied to the presented techniques: (1) Portability, (2) Model Complexity, and (3) Adaptability. These three criteria function as a tool for understanding how the human factors considerations of workload estimation for unknown tasks interact with the machine learning aspects of non-IID techniques, with the aim of determining each technique’s practical utility.